# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

### v2

@@ A.1 Multivariate data-driven stochastic models
$\begin{split}\Delta\mathbf{x}=\mathbf{x}(t+\Delta t)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))\Delta t+\mathbf{\Sigma}(\mathbf{x})\sqrt{\Delta t}\bm{\xi}(t)~{}\end{split}$ (14)
where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$. Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [6, 133, 7, 8]. The strategy is then to fit a rescaled version of Eq. 14 from observational data alone. Such rescaled equation takes the following form:
@@
      $\begin{split}\mathbf{x}(t+1)=\mathbf{x}(t)+\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}\bm{\xi}(t)~{}\end{split}$ (16)
@@
      $\begin{split}\mathbf{x}(t+1)=\mathbf{x}(t)+\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}(\mathbf{x}(t))\bm{\xi}(t)~{}\end{split}$ (17)
@@
A.2 Scalar data-driven stochastic models
@@ Acknowledgements.
Most of the ideas discussed in this work stem from collaborations and exchanges with colleagues and friends. I am grateful for the supportive and engaging environment of the Zanna Group at Courant, as well as for interactions within the broader M2Lines project. I thank Laure Zanna and Carlos Fernandez-Granda for their feedback during early stages of the project. I am especially grateful to Pasha Perezhogin, Rory Basinski, Chris Pedersen, and Andre Souza for their thoughtful feedback and for many enriching discussions on data-driven modeling of the Climate system. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.

### v3

@@ C.1 Multivariate data-driven stochastic models
[…]
where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$. Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [6, 136, 7, 8]. The strategy is then to fit a rescaled version of Eq. (18) from observational data alone. Such rescaled equation takes the following form:
@@
[…]
@@
[…]
@@
C.2 Scalar data-driven stochastic models
@@ Acknowledgements.
Most of the ideas discussed in this work stem from collaborations and exchanges with colleagues and friends. I am grateful for the supportive and engaging environment of the Zanna Group at Courant, as well as for interactions within the broader M2Lines project. I thank Laure Zanna and Carlos Fernandez-Granda for their feedback during early stages of the project. I am especially grateful to Pasha Perezhogin, Rory Basinski, Chris Pedersen, and Andre Souza for their thoughtful feedback and for many enriching discussions on data-driven modeling of the Climate system. I acknowledge an interesting email exchange with Prof. Tony Roberts, who drew my attention to recent advances on stochastic model reduction. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.